From GlueXWiki
Jump to: navigation, search

September 20, 2012 FDC meeting


  1. Production Construction Tracking (Dave)
  2. Cathode corrosion (Lubomir, Vlad)
    • Cathode protection tests
    • Recent SEM results [1], [2], [3], [4]
  3. Engineering (Bill)
  4. Electronics (Chris, Nick)
  5. Third package tests FDC E-log (Beni, Lubomir)
  6. TDR, pages 132-143 (Lubomir)
  7. Other


Participants: Bill, Dave, Glenn, Chris, Nick, Vlad, Simon, Beni, Eugene, and Lubomir


- Dave: we started working on two new cathodes, #72 and #73, using the last two sets of foils (new foils are coming November/December). Casey and Tina finished stringing today the new wire frame #29. There is a small delay with the new g10 wire frames (Vision Machine didn't have the vacuum pump), but they promised to have the first frame by Monday.

- The first three cells of package #1 were assembled Monday-Tuesday using the standard Viton+Apiezon technique. Since then it's being flushed with gas. Oxygen goes down very fast, now it's about 1000ppm and expect it to be ~300ppm tomorrow so Vlad will start testing the package.

Cathode corrosion

- We tried several techniques for covering the cathodes but all they had some problems:

  1. Kapton film + Hysol: many wrinkles, bubbles. Bill said he could do this using vacuum bagging, but should be done before the foils are stretched on the frame.
  2. Kapton film + Apiezon: the O-ring made a channel into the grease and the surface outside of that channel was elevated due to the grease covered with the film. The O-ring compression was low, several leaks were found especially between the films. Still the oxygen went down to 75ppm for a cell with a cathode half covered with this method.
  3. Hysol only: the epoxy layer that was applied was too thin (<0.5mil) so that it doesn't protect the copper efficiently. That was tested in the vacuum chamber were we saw copper discoloration with the Vition+Apiezon+Hysol sample similar to the sample without Hysol. It will be difficult to apply thicker Hysol layer evenly. The oxygen went down to 75ppm (combined test with the above method).
  4. Kapton tape with acrylic adhesive: Hera-B outer tracker had Molter effect and studied all the materials. The SCOTCH 467MP tape that they used showed out-gassing.

- Latest pictures from the SEM analyses are linked above. They correspond to a Viton+Apizon sample with the right pressure (using spacer ring) inside the vacuum chamber filled with gas. The above pictures/spectral analyses correspond to the dark area at one side of the O-ring. With 1K magnification only copper and carbon are seen and some small traces of sulfur. When focusing on the small white particle from the first picture, one sees much more sulfur. The conclusion could be that with Viton we have significantly less sulfur which is not enough to destroy the copper surface as this is the case with EPDM. It can just create tiny holes through the copper that we have seen before using HIREX microscope with back-light. Based on this and on the failure of the above methods (so far) we decided to use the standard Viton+Apiezon method. In principle one could improve the above technologies and do more tests, but it will take significant efforts and time without being sure about their success.

- Vlad will continue with the corrosion tests in the vacuum chamber at several high temperatures so that we can extrapolate the results to room temperatures.


- Bill: soon will start working on the cooling cart. He explained the FDC fiducial scheme and the survey procedures. There will be four targets on each package visible from the front/back even when the detector is installed and cabled. There will be a survey at Blue Crab of the packages as they are mounted on the cart, and also of the strong-back at the stringing table w.r.t the pins. The second survey will be done at the platform and the third one when the detector is installed.


- Nick and Chris populated one more PCB set to be used for wire frame #30. Nick gave Beni a LV module with distribution cards and cables enough for one package. The voltage at the LV source should be 6V to have at least 3V on each card (Nick). Eugene: one should measure the actual voltages at the cards and check if they will work when you remove one, two, or three pre-amps on the same cable.

Third package tests

- Lubomir showed pictures (entry 79 in FDC ELOG) from fADC125 to which one wire, one top strip and one down strip cards are connected, each 24 channels with same color on the plots. Shown is a "typical" event with one wire channel (blue) and several top (red) and bottom (green) strips fired. The second picture shows the noise on the strips, < 200 fADC units which is ~15mV, and much less on the wires.

- Beni showed (entry 78) results for the strip resolution perpendicular to the wire, by reconstructing the position of one wire, which is about ~400 microns that corresponds to the resolution along the wire of ~100 microns. Lubomir has done similar studies (entry 77) but using different charge integration on the strips and also for all the wires. The resolutions vary with the HV and are similar to the results with the full scale prototype. The main conclusion is that one can get resolutions below 150 microns.


- We are done with the draft TDR. Lubomir needs some numbers and re-formating of one picture from Bill. Eugene: maybe we should include some of the latest results and also explaining the algorithm used for the fADC data.